What pressure does a 3, point, 0, N,3.0N force on a 0, point, 60, c, m, squared,0.60cm
2
piston produce?
1
Expert's answer
2021-09-21T18:20:45-0400

"F = 3.0\,\mathrm{N}, \;\\ S = 0.60\,\mathrm{cm}^2 = 0.60\cdot10^{-4}\,\mathrm{m}^2 = 6\cdot10^{-5}\,\mathrm{m}^2"

We don't know the angle between the direction of the force and the normal to the piston. If the angle is "\alpha" then the pressure will be

"p = \dfrac{F\cos\alpha}{S}" , so the greatest pressure will correspond to "\alpha = 0", so the greatest pressure will be

"p = \dfrac{F}{S} = \dfrac{3.0\,\mathrm{N}}{6\cdot10^{-5}\,\mathrm{m}^2} = 5\cdot10^4\,\mathrm{Pa}."